### v2

# Reduction of Cosymplectic groupoids by cosymplectic moment maps

Abstract.
The Marsden-Weinstein-Meyer symplectic reduction has an analogous version for cosymplectic manifolds. In this paper we extend this cosymplectic reduction to the context of groupoids. Moreover, we prove how in the case of an algebroid associated to a cosymplectic groupoid, the integration commutes with the reduction (analogously to what happens in Poisson geometry). On the other hand, we show how the cosymplectic reduction of a groupoid induces a symplectic reduction on a canonical symplectic subgroupoid. Finally, we study what happens to the multiplicative Chern class associated with the $S^{1}$-central extensions of the reduced groupoid.
KEYWORDS. Group actions, cosymplectic groupoids, moment maps, co-symplectic reduction, Poisson manifolds.
1991 Mathematics Subject Classification:
53D20, 58H05, 58D19, 37J15, 53D17
Acknowledgments: We are grateful to David Iglesias for his valuable help in the discussions that inspired this work. We thank Cristian Ortiz for helpful and insightful discussions on the reduction of Poisson groupoids. We would like to thank Javier de Lucas and Xavier Rivas for inviting us to their Gamma seminar, where many ideas from this article were improved. López Garcia was supported by grant # 2022/04705-8, São Paulo Research Foundation (FAPESP).
1. Introduction
Cosymplectic manifolds arise as the odd counterpart of symplectic manifolds [22]. They can also be found as the underlying structures of other types of geometry as CR geometry [14], the odd-dimensional counterpart of Kähler manifolds and cokähler structures [2], almost contact manifolds [3], and in the description of time-dependent mechanics [19] and classical field theories [11]. Recently, these structures have been studied in the context of Lie groupoids giving rise to cosymplectic groupoids and their respective infinitesimal counterpart [13, 15] as central extensions of the underlying Poisson structure.
In the literature, several results about symmetries and reductions for geometric structures exist. In the case of cosymplectic geometry, we can consider two of them: the reduction of the associated Poisson structure and the Albert reduction using cosymplectic moment maps (analogous to the Marsden-Weinstein-Meyer symplectic reduction). The key difference between them is that the Poisson reduction does not preserve the nondegenerate condition in the cosymplectic case. On the other hand, the quotient of the level set of the cosymplectic moment map becomes cosymplectic again. For symplectic groupoids and Poisson structures on the units, these symmetries and reductions can be related by the integration theory of Lie groupoids [17]. The study of symmetries and reduction was later extended to Lie algebroids equipped with infinitesimal multiplicative 2-forms [6]. It is important to note that in [8], the case of IM 2-form reductions for quotients of vector bundles has been studied; in particular, Poisson structure quotients are studied. In contrast, our results focus on moment map reductions targeting co-symplectic structure reductions.
These notes extend the reduction technique in [6] to the particular case of cosymplectic groupoids and their Lie algebroids, taking into account the results in [15] and adapting the cosymplectic reduction theorem as in [1]. In Section 2 we start with an overview of cosymplectic manifolds, groupoids, and their infinitesimal data. In Section 3 we present the first main result of this article, which is a cosymplectic reduction by using a cosymplectic moment map in the context of Lie groupoids. Indeed, we consider symmetries of the cosymplectic forms equipped with a cosymplectic moment map that is also a groupoid morphism. Under certain conditions, the quotient of the 0-level set by the group action turns out to be a cosymplectic groupoid again (see Theorem 3.4).
Section 4 is intended to review some consequences of Theorem 3.4. In Section 4.1 we consider the algebroid of a cosymplectic groupoid and some symmetries at the infinitesimal level to define an algebroid morphism. This map can be used to perform a reduction at the infinitesimal level. On the other hand, using Lie’s second theorem we can integrate this algebroid morphism to a groupoid morphism which turns out to be a cosymplectic moment map. In Theorem 4.3, we show that the algebroid of the reduced cosymplectic groupoid corresponds to the reduction at the algebroid level.
Section 4.2 investigates how the cosymplectic reduction of a cosymplectic groupoid is related to the symplectic reduction of a particular symplectic subgroupoid. On a cosymplectic groupoid, there is a symplectic foliation on the space of arrows. The leaf containing the units is a symplectic subgroupoid. In Theorem 4.8, we show how cosymplectic reduction induces a Marsden-Weinstein symplectic reduction on the previously mentioned leaf. In section 4.3, using the relation between cosymplectic groupoids and $S^{1}$-central extensions with vanishing multiplicative Chern class, we study what happens to the multiplicative Chern class of the reduced groupoid. In Theorem 4.11, we consider an oversymplectic groupoid with an action of a compact group; if the multiplicative Chern class of the initial groupoid vanishes, then the same is true for the reduced groupoid.
Notation: For a groupoid $\mathcal{G}\rightrightarrows M$ we denote $t,s:\mathcal{G}\to M$ as target and source maps, $m:\mathcal{G}\times_{M}\mathcal{G}\to\mathcal{G}$ the product of composable arrows on the groupoid, and $\epsilon:M\to\mathcal{G}$ the units of $\mathcal{G}$. Throughout the article we assume that the space of arrows $\mathcal{G}$ is connected. Multiplicative forms $\theta$ in $\mathcal{G}$ are defined as those differential forms for which the equation $m^{*}\theta=pr_{1}^{*}\theta+pr_{2}^{*}\theta$ holds, where $pr_{j}:\mathcal{G}\times\mathcal{G}\to\mathcal{G}$ are the usual projection maps. In the case of an action of a Lie group $G$ on a manifold $Q$, for any $v\in\mathfrak{g}=Lie(G)$ we denote by $v_{Q}\in\mathfrak{X}(Q)$ the fundamental vector field corresponding to $v$.
2. Preliminaries
This section is dedicated to presenting the main objects of this article. Since we study the symmetries of cosymplectic groupoids and their infinitesimal counterpart, we present all the necessary definitions and tools to be used here..
2.1. Fundamental concepts about cosymplectic structures.
To establish the notation and as a quick reference we describe the relation between a cosymplectic manifold and its Poisson structure; for a more detailed description the reader can consult [15].
Let $(Q,\omega,\eta)$ be a cosymplectic manifold, i.e. $\omega\in\Omega_{cl}^{2}(Q)$ and $\eta\in\Omega_{cl}^{1}(Q)$, $\dim(Q)=2n+1$, and $\omega^{n}\wedge\eta$ defines a volume form. For this structure, the Lichnerowicz morphism $\flat:\mathfrak{X}(Q)\to\Omega^{1}(Q)$ given by
$\flat(X)=\imath_{X}\omega+(\imath_{X}\eta)\eta$ (1)
is an isomorphism. There is a remarkable vector field, called Reeb vector field, characterized by $\xi=\flat^{-1}(\eta)$, in particular it satisfies
$\imath_{\xi}\omega=0\qquad\mbox{\ and\ }\qquad\imath_{\xi}\eta=1.$ (2)
Example 2.1.
Let $(M,\omega)$ be a symplectic manifold of dimension $2m$. Let $i:N\hookrightarrow M$ be submanifold and $X$ be a symplectic vector field transverse to $N$. Then, $(N,i^{*}\omega,i^{*}(\imath_{X}\omega))$ is a cosymplectic manifold, in fact $d(i^{*}\omega)=0$, $di^{*}(\imath_{X}\omega))=i^{*}(\mathcal{L}_{X}\omega-\imath_{X}d\omega)=0$, and because $X$ is traverse to $N$, $\dim(N)=2(m-1)+1$ and $(i^{*}\omega)^{m-1}\wedge i^{*}(\imath_{X}\omega))$ is a volume form. Thus, the Lichnerowicz isomorphism is given by $\flat(Y)=i^{*}(\imath_{\tilde{Y}}\omega+(\imath_{\tilde{Y}}\imath_{X}\omega)\imath_{X}\omega),$ where $Y$ is $i$-related to $\tilde{Y}\in\mathfrak{X}(M)$.
Remark 2.2.
The converse is also true and was originally stated in [1, Prop 2]. The statement is as follows: if a manifold $Q$ is equipped with closed forms $\omega\in\Omega^{2}(Q)$ and $\eta\in\Omega^{1}(M)$ for which $\flat$ in (1) is an isomorphism and there exists a vector field as in (2), then $Q$ is odd-dimensional and $(Q,\omega,\eta)$ is a cosymplectic manifold. $\diamond$
Example 2.3.
Let $(M,\omega)$ be a symplectic manifold. Let $L$ be a 1-dimensional manifold with coordinate chart $r$. Then $M\times L$ with $\text{pr}_{1}^{*}\omega$ and $\eta=\text{pr}_{2}^{*}dr$ is a cosymplectic manifold. The Reeb vector field is $\xi:=\frac{\partial}{\partial r}$. Note that there is a regular symplectic foliation defined on $M\times L$, hence it is a Poisson manifold. If $f\in C^{\infty}(M\times L)$, then it is possible to define Hamiltonian vector fields on each leaf by $i_{X_{f_{r}}}\omega=df_{r}$, where $f_{r}=f(-,r)$. Furthermore, on $M\times L$,
$df_{(m,r)}=i_{X_{f_{r}}}\omega+\dfrac{\partial f}{\partial r}\text{pr}_{2}^{*}dr=i_{X_{f_{r}}}\omega+(\xi f)\eta$
is satisfied. Moreover, for $f,g\in C^{\infty}(M\times L)$ such that $i_{X_{g}}\eta=0$, the Poisson bracket is $\omega(X_{f},X_{g})$.
The particular case studied in Example 2.3 motivates the definition of cosymplectic Hamiltonian vector fields. Thus, given a smooth function $f$, the cosymplectic Hamiltonian vector field is defined by $X_{f}=\flat^{-1}(df-(\xi f)\eta)$, or equivalently by the unique vector field so that
$\imath_{X_{f}}\omega=df-\xi(f)\eta,\qquad\mbox{\ and\ }\qquad\imath_{X_{f}}\eta=0.$ (3)
Furthermore, the Poisson structure of Example 2.3 is a particular case of a more general result. It is known that $\ker(\eta)\subset TQ$ defines a regular integrable foliation $\mathcal{F}$ of corank 1 [15]. Moreover, the leaves of $\mathcal{F}$ are symplectic with respect to the restriction of $\omega$. Indeed, there is a one to one correspondence between cosymplectic structures on $Q$ and regular Poisson structures of corank 1 on $Q$ with a Poisson vector field transverse to the symplectic foliation (see [18]).
Hence, there is a Poisson structure $(Q,\pi_{Q})$, whose bracket satisfies $\{f,g\}=X_{f}g=\omega(X_{f},X_{g})$ because $i_{X_{g}}\eta=0$. As the Hamiltonian vector fields belong to $\mathrm{Ker}(\eta)$, we can verify that the bracket can also be written as $\{f,g\}=\omega(\flat^{-1}(df),\flat^{-1}(dg)).$ From the definition of $\xi$, we get $\mathcal{L}_{\xi}\omega=0$ and $\mathcal{L}_{\xi}\eta=0$. By using the Cartan formula and (3) we obtain
$\xi\{f,g\}=i_{X_{f}}i_{[\xi,X_{g}]}\omega-i_{X_{g}}i_{[\xi,X_{f}]}\omega=X_{f}(\xi g)-X_{g}(\xi f)=\{f,\xi g\}-\{g,\xi f\}.$
Therefore, the vector field $\xi$ is a Poisson vector field transverse to the leaves of $\mathcal{F}$.
A cosymplectic map $\phi:(Q,\omega,\eta)\to(Q^{\prime},\omega^{\prime},\eta^{\prime})$ is a map between cosymplectic manifolds such that $\phi^{*}\omega^{\prime}=\omega$ and $\phi^{*}\eta^{\prime}=\eta$, in particular $\xi$ and $\xi^{\prime}$ are $\phi$-related. If in addition $\phi$ is a diffeomorphism, then $i_{\phi_{*}X_{f}}\eta^{\prime}=0$ and
$i_{\phi_{*}X_{f}}\omega^{\prime}=(\phi^{-1})^{*}(i_{X_{f}}\omega)=d((\phi^{-1})^{*}f)-\xi^{\prime}((\phi^{-1})^{*}f)\eta^{\prime}.$
This relation says that $\phi_{*}X_{f}=X_{(\phi^{-1})^{*}f}$, and the map $\phi$ is a Poisson morphism between $\pi_{Q}$ and $\pi_{Q^{\prime}}$.
2.2. Symmetries and reduction
For this section, we assume a Lie group $G$ acting on a cosymplectic manifold on $(Q,\omega,\eta)$ by cosymplectic diffeomorphisms. Thus, it induces a $G$ Poisson action. In the case of a free and proper action, we obtain a reduced Poisson structure $\pi_{r}$ on $Q/G$ with quotient map $q:Q\to Q/G$ as Poisson morphism, that is $q_{*}\pi_{Q}=\pi_{r}$. By the correspondence between regular Poisson structures of corank 1 endowed with a Poisson vector field transverse to the symplectic foliation and cosymplectic structures, we have that a Poisson $Q/G$ reduction can be non-cosymplectic due to issues of rank or transversality. Nevertheless, there is a reduction procedure that preserve the cosymplectic structure, which is called cosymplectic reduction (see [1, 9]).
Let $G$ be a Lie group with Lie algebra $\mathfrak{g}=Lie(G)$. Let $(Q,\omega,\eta)$ be a cosymplectic manifold with an action of $G$ by cosymplectic morphisms. A smooth function $J:Q\to\mathfrak{g}^{*}$ is called a cosymplectic moment map if it is $G$-equivariant with respect to the coadjoint action, and for any $v\in\mathfrak{g}$ the Hamiltonian vector field $X_{\langle J,v\rangle}$ is the infinitesimal generator $v_{Q}\in\mathfrak{X}(Q)$. The value $0\in\mathfrak{g}^{*}$ is called a clean value if $J^{-1}(0)\subset Q$ is a submanifold and $\ker(dJ)_{x}=T_{x}J^{-1}(0)$ for all $x\in J^{-1}(0)$ (in [1, 23], it is called a weakly regular value). Clearly, a regular value is a clean value. Under these conditions there is a cosymplectic version of the classical Marsden-Weinstein reduction theorem [23]:
Theorem 2.4 ([1]).
Let $G$ be a Lie group acting on a cosymplectic manifold $(Q,\omega,\eta)$ by cosymplectic automorphisms. Let $J:Q\to\mathfrak{g}^{*}$ be a cosymplectic moment map such that $\xi(\langle J,v\rangle)=0$ for all $v\in\mathfrak{g}$. If 00 is a clean value of $J$, and $G$ acts freely and properly on $J^{-1}(0)$, then $Q_{red}=J^{-1}(0)/G$ is a cosymplectic manifold. The cosymplectic structure $(\omega_{red},\eta_{red})$ is defined by
$i^{*}\omega=p^{*}\omega_{red},\qquad i^{*}\eta=p^{*}\eta_{red},$
where $i:J^{-1}(0)\to Q$ is the inclusion and $p:J^{-1}(0)\to Q_{red}$ is the projection.
Remark 2.5.
Similarly to the symplectic case, when the action of $G$ on $Q$ is free, it follows that 0 is a regular value of $J$. From the Lichnerowicz isomorphism we observe that $v_{Q}=\flat^{-1}(d\langle J,v\rangle)$. Thus, $d\langle J_{x},v\rangle\neq 0$ if and only if $v_{Q}(x)\neq 0$ along $J^{-1}(0)$. Since $v_{Q}(e^{\tau v}x)=\left.\frac{d}{dt}\right|_{t=\tau}e^{tv}x$, then by the freeness of the action we conclude that $v_{Q}(x)\neq 0$ for any $v\neq 0\in\mathfrak{g}$. Consequently $dJ_{x}:T_{x}Q\to\mathfrak{g}^{*}$ is a surjective map. $\diamond$
Using a correspondence between cosymplectic moment maps on $Q$ and symplectic moment maps on $Q\times R$, we can study conditions under which cosymplectic moment maps exist.
Proposition 2.6.
Let $G$ be a Lie group acting on a connected cosymplectic manifold $(Q,\omega,\eta)$ by cosymplectic automorphisms such that $\imath_{v_{Q}}\eta=0,$ for all $v\in\mathfrak{g}$. If $H^{1}(Q)=0$ or $H^{1}_{Ch}(\mathfrak{g})=0$, then there exists a cosymplectic moment map $J:Q\to\mathfrak{g}^{*}$ such that $\xi(\langle J,v\rangle)=0$ for all $v\in\mathfrak{g}$.
Proof.
Let $i:Q\to Q\times\mathbb{R}$ the inclusion, and $p_{Q}$ and $p_{\mathbb{R}}$ be the canonical projections from $Q\times\mathbb{R}$ to $Q$ and $\mathbb{R}$, respectively. Consider $Q\times\mathbb{R}$ with the symplectic form $\tilde{\omega}=p_{Q}^{*}\omega+p_{\mathbb{R}}^{*}dt\wedge p_{Q}^{*}\eta$. There is a trivial extension of the $G$-action on $Q$ to a $G$-action on $Q\times\mathbb{R}$ given by $g\cdot(q,r)=(g\cdot q,r)$. We claim that there exists a one-to-one correspondence between cosymplectic moment maps on $(Q,\omega,\eta)$ satisfying $\xi\langle J,v\rangle=0$, and symplectic moment maps on $(Q\times\mathbb{R},\tilde{\omega})$.
Note that the vector fields $v_{Q}$ and $v_{Q\times\mathbb{R}}$ are $p_{Q}$-related. Thus, if $J:Q\to\mathfrak{g}^{*}$ is a cosymplectic moment map such that $\xi\langle J,v\rangle=0$, then $\tilde{J}:\mathbb{Q}\times\mathbb{R}\to\mathfrak{g}^{*}$ given by $\tilde{J}(q,r)=J(q)$ satisfies
$\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=\imath_{v_{Q\times\mathbb{R}}}(p_{Q}^{*}\omega+p_{\mathbb{R}}^{*}dt\wedge p_{Q}^{*}\eta)=p_{Q}^{*}(\imath_{v_{Q}}\omega)=p_{Q}^{*}d\langle J,v\rangle=d\langle\tilde{J},v\rangle.$
On the other hand, given a symplectic moment map $\tilde{J}:Q\times\mathbb{R}\to\mathfrak{g}^{*}$, we verify $\partial_{r}\langle\tilde{J},v\rangle=\imath_{\partial_{r}}\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=p_{Q}^{*}(\imath_{v_{Q}}\eta)=0.$ Thus, $\tilde{J}$ does not depend on $r$, and $J:=i^{*}\tilde{J}$ is well-defined. If $\tilde{\xi}=i_{*}(\xi)$, then $\imath_{\tilde{\xi}}\imath_{v_{Q\times\mathbb{R}}}\tilde{\omega}=0.$ Therefore, $\xi\langle J,v\rangle=\tilde{\xi}\langle\tilde{J},v\rangle=0.$ Consequently, from this correspondence we conclude that the obstruction to the existence of $\tilde{J}$ is the same as $J$. The result follows from [24, Prop. 4.5.17] and from the homotopy equivalence between $Q$ and $Q\times\mathbb{R}$. ∎
2.3. Cosymplectic groupoids and their infinitesimal counterpart.
The concept of cosymplectic groupoid was originally introduced in [13], and have recently been studied in [15]. This structure is defined as a Lie groupoid $\mathcal{G}\rightrightarrows M$ endowed with multiplicative differential forms $\omega\in\Omega_{cl}^{2}(\mathcal{G})$ and $\eta\in\Omega_{cl}^{1}(\mathcal{G})$, such that $(\mathcal{G},\omega,\eta)$ is a cosymplectic manifold.
Since the differential forms are multiplicative, the Poisson structure $(\mathcal{G},\pi_{\mathcal{G}})$ is multiplicative and the Reeb vector field $\xi$ is bi-invariant. In addition, it is possible to prove that the dimension of $\mathcal{G}$ is $2$dim$(M)+1$. On the other hand, let $\Sigma$ be the leaf of the symplectic foliation on $\mathcal{G}$ containing $M\hookrightarrow\mathcal{G}$. The groupoid $(\Sigma,\omega|_{\Sigma})\rightrightarrows M$ is symplectic and it integrates the Poisson structure $(M,\pi_{M})$, where $\pi_{M}=t_{*}\pi_{\mathcal{G}}$. For a detailed discussion, the reader may refer to [15].
Following the two canonical constructions of cosymplectic manifolds in Examples 2.1 and 2.3, as a corank 1 submanifold and the product by a 1-dimensional manifold, it is possible to produce examples of cosymplectic groupoids.
Example 2.7.
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a symplectic groupoid. Consider $i:\mathcal{G}_{0}\hookrightarrow\mathcal{G}$ a Lie subgroupoid over $M_{0}\hookrightarrow M$, both manifolds of codimension 1. If there exists a multiplicative symplectic vector field $X$ transverse to $\mathcal{G}_{0}$, then $(\mathcal{G}_{0},i^{*}\omega,i^{*}(\imath_{X}\omega))\rightrightarrows M$ is a cosymplectic groupoid. For a concrete example, consider a manifold $Q$, $\mathcal{G}=T^{*}(Q\times\mathbb{R})$ and $\omega=\omega_{can}$. Let $\mathcal{G}_{0}$ be the product of $T^{*}Q\rightrightarrows Q$ and ${}\mathbb{R}\rightrightarrows\{*\}$. If $r$ is the coordinate of the base of $T^{*}\mathbb{R}\to\mathbb{R}$, then the vector field $X=\partial r$ is multiplicative, symplectic and transverse to $\mathcal{G}_{0}$.
Example 2.8.
Associated with the construction of Example 2.3 we can consider two product groupoids. Let $(\mathcal{G}_{0},\omega)\rightrightarrows M$ be a symplectic groupoid. Consider $\mathcal{G}_{1}=\mathcal{G}_{0}\times S^{1}\rightrightarrows M$ and $\mathcal{G}_{2}=\mathcal{G}_{0}\times S^{1}\rightrightarrows M\times S^{1}$, where the structural maps of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ coming from the groupoids $S^{1}\rightrightarrows\{*\}$ and $S^{1}\rightrightarrows S^{1}$, respectively. The cosymplectic forms $pr_{\mathcal{G}_{0}}^{*}\omega$ and $pr_{S^{1}}d\theta$ are multiplicative in $\mathcal{G}_{1}$ making it a cosymplectic groupoid. However, $(\mathcal{G}_{2},\text{pr}_{\mathcal{G}_{0}}^{*}\omega,\text{pr}_{S^{1}}^{*}d\theta)$ is not a cosymplectic groupoids by dimensional issues.
We now proceed to describe the infinitesimal counterpart of these groupoids. For this, we recall the theory of infinitesimal multiplicative forms, as in [5]. These structures are defined on the Lie algebroid $A=Lie(\mathcal{G})$. For any multiplicative form $\alpha\in\Omega_{cl}^{k}(\mathcal{G})$ we define the vector-bundle map $\mu:A\to\Lambda^{k-1}T^{*}M$ covering the identity, defined by
$\mu:A\to\Lambda^{k-1}T^{*}M\qquad\langle\mu(u),X_{1},\ldots,X_{k-1}\rangle:=\alpha(u,d\epsilon X_{1},\ldots,d\epsilon X_{k-1}),$
where $u\in A$ and $X_{i}\in TM$. This definition can also be written as $\epsilon^{*}(\imath_{u^{R}}\alpha)=\mu(u)$, where $u^{R}$ is the right invariant vector field on $\mathcal{G}$ induced by $u\in A$. The bundle map $\mu$ is called IM k-form and it satisfies the compatibility conditions
$\imath_{\rho(a)}\mu(b)$ $=-\imath_{\rho(b)}\mu(a),$ (4)
$\mu([a,b])$ $=\mathcal{L}_{\rho(a)}\mu(b)-\imath_{\rho(b)}d\mu(a),$ (5)
for any $a,b\in\Gamma(A)$. For a cosymplectic groupoid $(\mathcal{G},\omega,\eta)$, $\omega$ induces a IM 2-form $\mu:A\to T^{*}M$ satisfying previous two conditions, while the multiplicative form $\eta$ has an associated IM 1-form $\nu:A\to\mathbb{R}$ satisfying the relation
$\nu([a,b])=\rho(a)\nu(b)-\rho(b)\nu(a).$ (6)
Since $\omega$ and $\eta$ are multiplicative, the map $a\to(\mu(a),\nu(a))$ between the Lie algebroid $A$ and the central extension algebroid $T^{\ast}M\oplus\mathbb{R}$ is an isomorphism. Under this identification, the anchor map and bracket are
$\rho(df,\lambda)=X_{f},\qquad[(df_{1},g_{1}),(df_{2},g_{2})]=(d\{f_{1},f_{2}\}_{M},X_{f_{1}}(g_{2})-X_{f_{2}}(g_{1})),$ (7)
where the Hamiltonian vector fields are defined by the Poisson structure $(M,\pi_{M})$. For more details on these structures, the reader may refer to [15, Section 3]. Note that in order to define the central extension algebroid $T^{\ast}M\oplus\mathbb{R}$, it is only necessary that $M$ has a Poisson structure.
3. Symmetries and reduction of cosymplectic groupoids
We present here our first main contribution, which is a reduction theorem for the cosymplectic groupoid using a Marsden-Weinstein approach. We begin by pointing out why a Poisson reduction is not convenient when working with cosymplectic groupids. This motivates the introduction of a moment map in cosymplectic groupids and thus using Theorem 2.4 to carry out the reduction.
[…]
${(\mathcal{G},\pi_{\mathcal{G}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t}$${(\mathcal{G}/G,\pi_{r})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{t_{r}}$${(M,\pi_{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(M/G,\pi_{r}).}$
It is important to highlight that this reduction may not be a cosymplectic groupoid. In the following example, we start with a cosymplectic groupoid and after performing a Poisson reduction we obtain a groupoid that is not a regular Poisson groupoid of corank 1, and consequently it is not a cosymplectic groupoid.
Example 3.1.
Consider the cosymplectic groupoid $\mathcal{G}=T^{*}\mathbb{R}^{n}\times S^{1}\rightrightarrows\mathbb{R}^{n}$ of Example 2.8, with $\omega=\omega_{can}$. Let $\mathbb{R}^{n-k}\curvearrowright\mathbb{R}^{n}$ be the translation in the last $n-k$ coordinates, and $\varphi$ the $\mathbb{R}^{n-k}$-action on $T^{*}\mathbb{R}^{n}\times S^{1}$ defined by $\varphi_{\alpha}((q,p),\beta)=((\alpha\cdot q,p),\beta)$. Since $\mathcal{G}$ is a $\mathbb{R}^{n-k}$-invariant cosymplectic groupoid, then we obtain the Poisson groupoid $\mathcal{G}/\mathbb{R}^{n-k}\rightrightarrows\mathbb{R}^{n}/\mathbb{R}^{n-k}$. Note that $\dim(\mathcal{G}/\mathbb{R}^{n-k})=n+k+1$ and $\dim(\mathbb{R}^{n}/\mathbb{R}^{n-k})=k$, thus for $k\neq n$ the reduction is not a cosymplectic groupoid.
Thus, given a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ and a Lie group $G$ acting on $\mathcal{G}$ by groupoid automorphisms and cosymplectic diffeomorphisms, if we want to study cosymplectic reduction, we must consider a moment map. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map. In addition, we assume that $\mathcal{J}$ is a morphism of Lie groupoids, where $\mathfrak{g}^{*}$ is considered as an Abelian group with respect to addition. If we suppose that 0 is a clean value for $\mathcal{J}$, then it is possible to show that $\mathcal{J}^{-1}(0)\subset\mathcal{G}$ is a wide Lie subgroupoid. The following result is strongly motivated in [6].
Lemma 3.2.
Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a morphism of groupoids, where $\mathfrak{g}^{*}$ is considered as an Abelian group with respect to addition. If 00 is a clean value of $\mathcal{J}$, then $\mathcal{J}^{-1}(0)\subset\mathcal{G}$ is a wide Lie subgroupoid.
Proof.
For a pair of composable arrows $(g,h)$, we have $\mathcal{J}(gh)=\mathcal{J}(g)+\mathcal{J}(h)$. Hence, $\mathcal{J}^{-1}(0)$ is closed under the multiplication and inverse on $\mathcal{G}$. Also, for any $x\in M$, $\epsilon(x)\in\mathcal{J}^{-1}(0)$, so the subgroupoid is wide. What remains to be demonstrated is that the restrictions of the maps $s,t$ to $\mathcal{J}^{-1}(0)$ are submersions. If $i:\mathcal{J}^{-1}(0)\to\mathcal{G}$ is the inclusion map, then
$\ker(d(i^{*}s))_{g}=\ker(ds)_{g}\cap T_{g}\mathcal{J}^{-1}(0)=\ker(ds)_{g}\cap\ker(d\mathcal{J})_{g},\quad\text{ for all }g\in\mathcal{G}.$
Thus, $\text{rank}(d(i^{*}s))=\dim(M)$. ∎
It is worth to remark that since $G$ acts freely and properly on $M$, then $G$ also acts freely and properly on $\mathcal{G}$ (see [17, Prop. 4.4]). A direct consequence of this fact is that it is enough to consider a free action on the units of the groupoid to verify that 0 is a regular value of $\mathcal{J}$.
Lemma 3.3.
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid automorphisms. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map, and $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$. If the action $G$ acts freely on $M$, then 00 is a regular value of $\mathcal{J}$.
Proof.
Since the action is free in $M$, then it is also free in $\mathcal{G}$ (see [17]). Thus we conclude the result using Remark 2.5. ∎
Motivated by Theorem 2.4, we propose to address the cosymplectic structure of $\mathcal{J}^{-1}(0)/G$, and see if it has a compatible groupoid structure. Thus, we present the first main theorem of these notes, which is a Marsden–Weinstein reduction of cosymplectic groupoids.
Theorem 3.4.
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid automorphisms. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a cosymplectic moment map such that it is a Lie groupoid morphism, and $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$. If $G$ acts freely and properly on $M$, then $\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G$ is a cosymplectic groupoid over $M/G$. The cosymplectic structure $(\omega_{red},\eta_{red})$ is defined by
$i^{*}\omega=P^{*}\omega_{red},\qquad i^{*}\eta=P^{*}\eta_{red},$
where $i:\mathcal{J}^{-1}(0)\to\mathcal{G}$ is the inclusion and $P:\mathcal{J}^{-1}(0)\to\mathcal{G}_{red}$ is the projection.
Proof.
From Lemma 3.3, it follows that $\mathcal{J}^{-1}(0)$ is a smooth manifold. Since $G$ acts freely and properly on $\mathcal{J}^{-1}(0)$, the $G$-equivariance gives rise to a smooth manifold $\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G$. Analogously, $M/G$ is a smooth manifold. Provided that the structural maps of the groupoid $\mathcal{G}\rightrightarrows M$ are $G$-equivariant, then they descend to $\mathcal{G}_{red}$. Thus $\mathcal{G}_{red}$ has a natural groupoid structure on $M/G$. Let $P:\mathcal{J}^{-1}(0)\to\mathcal{G}_{red}$ and $p:M\to M/G$ be the quotient maps. They are surjective submersions, and $P$ covers $p$. Furthermore, the pair $(P,p)$ defines a groupoid morphism between $\mathcal{J}^{-1}(0)$ and $\mathcal{G}_{red}$. Thus, the source and target map on $\mathcal{G}_{red}$ are submersions.
Moreover, by Theorem 2.4, the manifold $(\mathcal{G}_{red},\omega_{red},\eta_{red})$ characterized by $i^{*}\omega=P^{*}\omega_{red}$ and $i^{*}\eta=P^{*}\eta_{red},$ is a cosymplectic manifold. To show that $\mathcal{G}_{red}\rightrightarrows M/G$ is a cosymplectic groupoid remains to verify that $\omega_{red}$ and $\eta_{red}$ are multiplicative. Since the quotient map commutes with the structural groupoids maps, and by Lemma 3.2 we have the next commutative diagram
${\mathcal{J}^{-1}(0)^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{0}}$$\scriptstyle{(i,i)}$$\scriptstyle{(P,P)}$${\mathcal{J}^{-1}(0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$$\scriptstyle{P}$${\mathcal{G}_{red}^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m_{red}\hskip 14.22636pt}$${\mathcal{G}_{red}}$${\mathcal{G}^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{m}$${\mathcal{G}.}$
From which we conclude,
$P^{*}m_{red}^{*}\eta_{red}=m_{0}^{*}i^{*}\eta=i^{*}(pr_{1}^{*}\eta+pr_{2}^{*}\eta)=P^{*}(pr_{1}^{*}\eta_{red}+pr_{2}^{*}\eta_{red}),$
where $m_{0}$ is the restriction of the product in $\mathcal{G}$ to $\mathcal{J}^{-1}(0)$. Because $P$ is a surjective submersion, then $m_{red}^{*}\eta_{red}=pr_{1}^{*}\eta_{red}+pr_{2}^{*}\eta_{red}$. Analogously, it is proved that $\omega_{red}$ is multiplicative. ∎
Example 3.5.
In Example 2.7, consider a Lie group $G$ acting on $(\mathcal{G},\omega)$ by symplectic groupoid automorphisms. Let $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ be a symplectic moment map such that it is a Lie groupoid morphism. Suppose that $\mathcal{G}_{0}$ and $M_{0}$ are closed under the action of $G$. In addition, we assume that the vector field $X\in\mathfrak{X}(\mathcal{G})$ is $G$-invariant, and $X\langle\mathcal{J},v\rangle=0$ for all $v\in\mathfrak{g}$. Therefore, the map $\mathcal{J}_{0}:\mathcal{G}_{0}\xrightarrow{i_{0}}\mathcal{G}\xrightarrow{\mathcal{J}}\mathfrak{g}^{*}$ is a cosymplectic moment map and a Lie groupoid morphism. If $G$ acts freely and properly on $M$, and $G$ acts properly on $M_{0}$, then $(\mathcal{G}_{0})_{red}=\mathcal{J}_{0}^{-1}(0)/G\rightrightarrows M_{0}/G$ is a cosymplectic groupoid. Giving rise to the commutative diagram
${\mathcal{J}_{0}^{-1}(0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{J}^{-1}(0)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{G}_{0}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{0}\hskip 17.07164pt}$${\mathcal{G}}$${(\mathcal{G}_{0})_{red}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${\mathcal{G}_{red},}$
where $(\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G,\omega_{red})\rightrightarrows M/G$ is the reduced symplectic groupoid [17]. Moreover, there exists a $G$-invariant vector field $X_{0}\in\mathfrak{X}(\mathcal{J}^{-1}(0))$, which is $i$-related to $X$. Also, $X_{0}$ is projectable to a vector field $\bar{X}_{0}\in\mathfrak{X}(\mathcal{G}_{red})$. Thus, the reduced cosymplectic groupoid coincides with $((\mathcal{G}_{0})_{red},i^{*}\omega_{red},i^{*}(\imath_{\bar{X_{0}}}\omega_{red}))$.
Example 3.6.
In Example 3.1, consider the linear momentum associated with $\mathbb{R}^{n-k}\curvearrowright T^{*}\mathbb{R}^{n}$. Thus, the cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathbb{R}^{n-k}$ is defined by $\mathcal{J}((q,p),\beta)=(p_{k+1},\ldots,p_{n})$. Then, $\mathcal{J}^{-1}(0)=T^{*}\mathbb{R}^{k}\times\mathbb{R}^{n-k}\times S^{1}$ and $\mathcal{G}_{red}=T^{*}\mathbb{R}^{k}\times S^{1}$. Note that the inclusion $i:\mathcal{J}^{-1}(0)\hookrightarrow T^{*}\mathbb{R}^{n}\times S^{1}$ is by considering the zero section in the last $n-k$ coordinates, thus $i^{*}\omega_{can}=\sum_{j=1}^{k}dq_{j}\wedge dp_{j}$. Therefore $\omega_{red}=pr_{T^{*}\mathbb{R}^{k}}^{*}\sum_{j=1}^{k}dq_{j}\wedge dp_{j}$ and $\eta_{red}=pr_{S^{1}}d\theta$.
4. Consequences of the main theorem
This section is devoted to applying Theorem 3.4 to specific situations concerning the multiplicativity of cosymplectic forms in a Lie groupoid. In particular, we are inspired by the methods presented in [6, 17] but in the context of cosymplectic structures. We also explore some of the main results in [15] when symmetries and reduction are considered.
4.1. Symmetries and reduction at the infinitesimal level
Under the existence of symmetries of a Lie algebroid of a cosymplectic groupoid, it is interesting to study how integration is related to reduction. Motivated by the results in [6], we study the reductions of the IM-forms associated with the cosymplectic forms.
We state the geometrical setup as follows. Let $M$ be a Poisson manifold. Consider a Lie algebroid $A\Rightarrow M$, endowed with an IM 2-form and an IM 1-form $(\mu,\nu)$ so that $A\cong T^{*}M\oplus\mathbb{R}$ as in (4)-(7). In addition, let $G$ be a Lie group acting on the manifold $M$. Suppose that this action lifts to an action on $A\Rightarrow M$ by Lie algebroid automorphisms, such that $\mu$ is $G$-equivariant and $\nu$ is $G$-invariant, i.e.
$\mu\circ\varphi_{g}^{A}=(d\varphi_{g^{-1}}^{M})^{*}\circ\mu,\qquad\nu\circ\varphi_{g}^{A}=\nu$
for all $g\in G$. The superindex notation refers to the space where $G$ acts.
Let $J_{can}:T^{*}M\to\mathfrak{g}^{*}$ be the canonical symplectic moment map of the cotangent lifting of the $G$-action on $M$ to $T^{*}M$. It is defined by $\langle J_{can}(\alpha),v\rangle=\alpha(v_{M})$ for any $v\in\mathfrak{g}$. Thus, we obtain a Lie algebroid morphism
[…]
where $\mathfrak{g}^{*}$ is thought as a Lie algebroid over a point with trivial Lie bracket. Note that, this map is $G$-equivariant because maps $\mu$ and $\mathcal{J}_{can}$ are $G$-equivariant. Moreover, $J$ is a Lie algebroid morphism (see [6]).
From now on we assume that $J^{-1}(0)\subset A$ is a subbundle. If $\mu$ has constant rank, the former condition is equivalent to $\mu(A)\cap J_{can}^{-1}(0)$ having constant rank. In [6], the authors prove that the subbundle $J^{-1}(0)\subset A$ is also a Lie subalgebroid. Furthermore, they show that the quotient by the action $G$ inherits a unique Lie algebroid structure. If the action $G\curvearrowright M$ is free and proper, then $M/G$ is a smooth manifold with surjective submersion quotient map $p:M\to M/G$. In addition, since $J^{-1}(0)$ is $G$-invariant, we can consider the vector bundle $A_{red}\rightarrow M/G$. It is defined as the unique vector bundle over $M/G$ such that the pullback bundle $p^{*}(A_{red})$ is naturally isomorphic to $J^{-1}(0)$, and the pullback of sections defines an identification of $\Gamma(A_{red})$ with the space of $G$-invariant sections of $\Gamma(J^{-1}(0)).$
Thus, the vector bundle $A_{red}$ inherits a unique Lie algebroid structure $(A_{red},[\cdot,\cdot]_{red},\rho_{red})$ so that:
  • The pullback of sections $p^{*}:\Gamma(A_{red})\to\Gamma(J^{-1}(0))$ preserves the Lie brackets.
  • For any $a\in\Gamma(A_{red})$ the vector field $\rho_{red}(a)$ and $\rho(p^{*}a)$ are $p$-related.
Let $\mu_{0}:J^{-1}(0)\to J_{can}^{-1}(0)$ and $\nu_{0}:J^{-1}(0)\to\mathbb{R}$ be the restrictions of the IM-forms by the inclusion map $J^{-1}(0)\to A$. Note that $\mu_{0}$ and $\nu_{0}$ are $G$-equivariant and $G$-invariant, respectively. Hence, it is possible to define the bundle maps $\mu_{red}:A_{red}\to T^{*}(M/G)$ and $\nu_{red}:A_{red}\to\mathbb{R}$ by satisfying the identities
$p^{*}(\mu_{red}(a))=\mu_{0}(p^{*}(a))\quad\text{ and }\quad p^{*}(\nu_{red}(a))=\nu_{0}(p^{*}(a)),$ (9)
where $a\in\Gamma(A_{red})$. The following diagrams illustrate the pullback maps that appear in these equations,
[…]
As presented below, a straightforward computation shows that the bundle maps $\mu_{red}$ and $\nu_{red}$ satisfy equations (4), (5), and (6), respectively.
Lemma 4.1.
The forms $\mu_{red}$ and $\nu_{red}$ are IM-form on $A_{red}$.
Proof.
The skew-symmetry of $\mu_{red}$ follows directly from the definition. To check bracket compatibility, consider $a,b\in\Gamma(A_{red})$, and note that
$p^{*}\mu_{red}([a,b])$ $=\mu_{0}([p^{*}a,p^{*}b])=\mathcal{L}_{\rho(p^{*}a)}\mu_{0}(p^{*}b)-\imath_{\rho(p^{*}b)}d\mu_{0}(p^{*}a)$
$=\mathcal{L}_{\rho(p^{*}a)}p^{*}\mu_{red}(b)-\imath_{\rho(p^{*}b)}dp^{*}\mu_{red}(a)=p^{*}(\mathcal{L}_{\rho_{red}(a)}\mu_{red}(b)-\imath_{\rho_{red}(b)}d\mu_{red}(a)),$
and
$p^{*}\nu_{red}([a,b])$ $=\nu_{0}([p^{*}a,p^{*}b])=\rho(p^{*}a)\nu_{0}(p^{*}b)-\rho(p^{*}b)\nu_{0}(p^{*}a)$
$=\rho(p^{*}a)p^{*}\nu_{red}(b)-\rho(p^{*}b)p^{*}\nu_{red}(a)=p^{*}(\rho_{red}(a)\nu_{red}(b)-\rho_{red}(n)\nu_{red}(a)).$
We now assume that $A$ is integrable by a source simply-connected groupoid. Let $(\mathcal{G},\omega,\eta)\rightrightarrows M$ be a cosymplectic groupoid, such that $Lie(\mathcal{G})=A$. The IM-forms $(\mu,\nu)$ defined by the forms $(\omega,\eta)$ satisfy (4)-(7) (see [15]). Let $G\curvearrowright A$ be an action as defined at the beginning of this section. From the Lie’s second theorem, the action $G\curvearrowright A$ integrates to an action $G\curvearrowright\mathcal{G}$ by Lie groupoid automorphisms. Notice that these two actions cover the $G$-action on $M$.
The Lie algebroid morphism $J:A\to\mathfrak{g}^{*}$ integrates to a $G$-equivariant Lie groupoid morphism
$\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*},$ (10)
Furthermore, it satisfies $\imath_{v_{\mathcal{G}}}\omega=-d\langle\mathcal{J},v\rangle$ for all $v\in\mathfrak{g}$ (see [6, Prop. 3.2]). Since $\mathcal{G}$ is a cosymplectic groupoid, the map $\mathcal{J}$ becomes a cosymplectic momentum map, as shown below.
Lemma 4.2.
The action of $G$ on $\mathcal{G}$ is by cosymplectic diffeomorphism. Moreover, the map $-\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ defined in (10) is a cosymplectic moment map, and $\xi\langle\mathcal{J},v\rangle=0$ for all $v\in\mathfrak{g}$.
Proof.
To verify that $\omega$ is $G$-invariant, we use that that $\mu$ is $G$-equivariant. Note that
$(\varphi_{g}^{\mathcal{G}})^{*}\omega(u,d\epsilon X)$ $=\omega(d\varphi_{g}^{\mathcal{G}}(u),d\varphi_{g}^{\mathcal{G}}(d\epsilon X))=\omega(d\varphi_{g}^{\mathcal{G}}(u),d\epsilon(d\varphi_{g}^{M}X))$
$=(d\varphi_{g^{-1}}^{M})^{*}\mu(u)(d\varphi_{g}^{M}X)=\mu(u)(X)=\omega(u,d\epsilon X),$
where $u\in A$ and $X\in TM$. Thus, from [5, Thm. 4.6] we conclude that $(\varphi_{g}^{\mathcal{G}})^{*}\omega=\omega$. Similarly, $G$-invariance of $\eta$ is proved by using that $\nu$ is $G$-invariant. To prove that $-\mathcal{J}$ is cosymplectic moment map we use that the $G$-action on $\mathcal{G}$ covers the action on $M$. In particular, we have $v_{\mathcal{G}}(x)=v_{M}(x)\in T_{x}M$ for $v\in\mathfrak{g}$ and $x\in M$. Because $\eta$ is multiplicative, then $\epsilon^{*}\eta=0$. Thus, $TM\subset\ker\eta$ and $(\imath_{v_{\mathcal{G}}}\eta)|_{M}=0$. Since $G$ preserves the form $\eta$, we deduce $0=\mathcal{L}_{v_{\mathcal{G}}}\eta=d\imath_{v_{\mathcal{G}}}\eta$. Therefore, $\imath_{v_{\mathcal{G}}}\eta=0.$ Furthermore, from $\imath_{v_{\mathcal{G}}}\omega=-d\langle\mathcal{J},v\rangle$, we conclude $-\xi\langle\mathcal{J},v\rangle=\imath_{\xi}\imath_{v_{\mathcal{G}}}\omega=0$. ∎
If $G$ acts freely on $M$, then by Lemma 3.3, 0 is a regular value of $\mathcal{J}$. Let $i:\mathcal{J}^{-1}(0)\to\mathcal{G}$ the inclusion map. Thus, the restricted source map satisfies $\ker(d(i^{*}s))_{g}=\ker(ds)_{g}\cap\ker(d\mathcal{J})_{g}$. Hence, we verify
$\ker(d(i^{*}s))|_{M}=A\cap\ker(d\mathcal{J})|_{M}=J^{-1}(0),$
making $\mathcal{J}^{-1}(0)$ an integration of $J^{-1}(0)$. If $G$ also acts properly on $M$, the quotient $\mathcal{G}_{red}=\mathcal{J}^{-1}(0)/G$ is a smooth manifold. We further recall that the action $G\curvearrowright\mathcal{G}$ is by groupoid automorphisms, so the quotient map $P:\mathcal{J}^{-1}(0)\to\mathcal{G}_{red}$ is a groupoid morphism. By differentiating $P$, we get $dP|_{J^{-1}(0)}:A_{red}\to Lie(\mathcal{G}_{red}).$ The kernel of $dP|_{J^{-1}(0)}$ is trivial, thus there is an identification between $A_{red}$ and $Lie(\mathcal{G}_{red})$ (see [6, Prop. 5.2]).
We summarize the results of this section in the following theorem, which is the main result of this section. In the next theorem, we suppose the source simply connected groupoid $\mathcal{G}$.
Theorem 4.3.
Let $(\mathcal{G},\omega,\eta)\rightrightarrows M$ be a cosymplectic groupoid and $(A,\mu,\nu)$ its Lie algebroid. Let $G$ be a Lie group acting freely and properly on $M$. Suppose that this action lifts to an action on $A$ by Lie algebroid automorphisms, such that $\mu$ is $G$-equivariant and $\nu$ is $G$-invariant. Let $J$ be the map in (8), and $\mathcal{J}$ the map in (10). If $J^{-1}(0)\subset A$ is a subbundle, then the reduced Lie algebroid $(A_{red},\mu_{red},\nu_{red})$ is integrable by the cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$.
Proof.
The algebroid $(A,\mu,\nu)$ satisfies the conditions of the geometrical setup established at the beginning of this section, where the Poisson structure on $M$ is given by $\pi_{M}=t_{*}\pi_{\mathcal{G}}$. By using the Lie algebroid morphism $J:A\to\mathfrak{g}^{*}$ defined in (8), we consider the algebroid $(A_{red},\mu_{red},\nu_{red})$ over $M/G$. Furthermore, the forms $\mu_{red}$ and $\nu_{red}$ are IM-forms by Lemma 4.1.
By Lie’s second theorem we integrate the $G$-action on $A$ to a $G$-action on $\mathcal{G}$, and we also integrate the map $J:A\to\mathfrak{g}^{*}$ to a groupoid morphism $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$. Lemma 4.2 states that the map $\mathcal{J}$ is also a cosymplectic moment map. By using Theorem 3.4, we conclude that $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ is a cosymplectic Lie groupoid. As mentioned above, $A_{red}=Lie(\mathcal{G}_{red})$. Furthermore, let $u\in A_{red}$, then using the groupoid morphism $(P,p):(\mathcal{J}^{-1}(0)\to M)\to(\mathcal{G}_{red}\to M/G)$, we have
$p^{*}\epsilon^{*}(\imath_{u^{R}}\eta_{red})$ $=\epsilon^{*}P^{*}(\imath_{u^{R}}\eta_{red})=\epsilon^{*}(\imath_{p^{*}(u^{R})}P^{*}\eta_{red})=\epsilon^{*}(\imath_{p^{*}(u^{R})}i^{*}\eta)$
$=\nu_{0}(p^{*}(u))=p^{*}(\nu_{red}(u)).$
Since $p$ is a surjective submersion we get $\epsilon^{*}(\imath_{u^{R}}\eta_{red})=\nu_{red}(u)$. Analogously for $\mu_{red}$. ∎
In other words, for a cosymplectic groupoid the reduction of its infinitesimal counterpart is the same as the infinitesimal version of its cosymplectic reduction. Thus, we have that
${(\mathcal{G},\omega,\pi)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$Liered${(\mathcal{G}_{red},\omega_{red},\pi_{red})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$Lie${(A,\mu,\nu)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(A_{red},\mu_{red},\nu_{red}),}$
commutes. The following is an immediate consequence of this theorem.
Corollary 4.4.
Let $(T^{*}M\oplus\mathbb{R},\mu=pr_{1},\nu=pr_{2})$ be a central extension Lie algebroid with brackets as in (7). If there is a $G$ action on $M$ so that lifts to action by Lie algebroid morphism on $A$ so that $\mu$ is equivariant and $\nu$ is invariant, then
$(T^{*}(M/G)\oplus\mathbb{R},\mu_{red}=pr_{1},\nu_{red}=pr_{2})$
is also a central extension algebroid isomorphic to $A_{red}$.
4.2. Reduction of the symplectic leaf
A natural question that arises is the symplectic reduction of the leaves of a cosymplectic groupoid. In this section we study how symplectic and cosymplectic reduction are related in the restriction to the symplectic leaf containing the units of a cosymplectic groupoid.
Given a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, let $\Sigma$ be the symplectic leaf of $\pi_{\mathcal{G}}$ containing $M$. Let $L:\Sigma\to\mathcal{G}$ denote the inclusion. In [15] it is proved that $\Sigma\subset\mathcal{G}$ is a Lie subgroupoid, and $(\Sigma,L^{*}\omega)\rightrightarrows M$ is a symplectic groupoid integrating $\pi_{M}$. Assuming that we impose symmetries on the cosymplectic groupoid $(\mathcal{G},\omega,\eta)$ so that we can construct its cosymplectic reduced groupoid, we inquire about the symmetries and the reduction of $(\Sigma,L^{*}\omega)\rightrightarrows M$. We want to restrict the action $G$ to the leaf $\Sigma$, and we want to define a symplectic moment map.
To show that the action of $G$ on $\mathcal{G}$ is restricted to an action on $\Sigma$, we first state a basic fact about morphisms between Poisson varieties.
Lemma 4.5.
Let $M$ and $N$ be Poisson manifold, and $\varphi:M\to N$ a Poisson diffeomorphism. Then, $\varphi$ preserves the symplectic foliation.
Proof.
Because $\varphi$ is a Poisson morphism, the Hamiltonian vectors are $\varphi$-related. Since the symplectic foliations in $M$ and $N$ are generated by their respective Hamiltonian vector fields, then the leaves in $M$ correspond to the leaves in $N$. In fact, let $f\in C^{\infty}(M)$ and $p,p^{\prime}\in M$ be on a same leaf. Let $\gamma$ be an integral curve of $X_{f}\in\mathfrak{X}(M)$ joining $p$ and $p^{\prime}$, therefore the curve $\tilde{\gamma}=\varphi\circ\gamma$ is the solution of $\varphi_{*}X_{f}\in\mathfrak{X}(N)$ joining $\varphi(p)$ and $\varphi(p^{\prime})$. ∎
It is important to highlight that the leaves of the symplectic foliation may be non-invariant by the action of a group by Poisson automorphisms (cf. [24, Ex. 4.2.5]). Moreover, in the case of standard momentum maps on Poisson manifolds, a necessary condition for the existence of a momentum map is that the action by Poisson morphisms being leaf preserving, which does not happen in general [24, §4.5]. However, in the case of a cosymplectic groupoid, the action does preserve the symplectic leaf containing the units. Thus, we can restrict the action to this leaf.
Lemma 4.6.
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$ by cosymplectic groupoid automorphisms. Let $\Sigma$ be the symplectic leaf of $\pi_{\mathcal{G}}$ containing $M$. Then, we can restrict the action $G\curvearrowright\mathcal{G}$ to a $G$-action on $\Sigma$, preserving the form $L^{*}\omega$.
Proof.
The action $G\curvearrowright\mathcal{G}$ is by Poisson diffeomorphisms. By Lemma 4.5, any left multiplication either transforms the leaf $\Sigma$ into another leaf of the symplectic foliation, or leaves it invariant. Furthermore, for any $g\in G$, the units $M$ are contained in $\varphi_{g}(\Sigma)$. Therefore, by the definition of $\Sigma$, we have $\varphi_{g}(\Sigma)=\Sigma$. In addition we verify that
$(\varphi^{\Sigma}_{g})^{*}L^{*}\omega=L^{*}(\varphi_{g}^{\mathcal{G}})^{*}\omega=L^{*}\omega.$
[…]
Lemma 4.7.
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ a cosymplectic moment map satisfying the conditions of Theorem 3.4. Let $\Sigma$ the symplectic leaf containing $M$. Then the groupoid morphism $\mathcal{J}_{0}:(\Sigma,L^{*}\omega)\to\mathfrak{g}^{*}$ defined by $L^{*}\mathcal{J}$ is a symplectic moment map with respect to the $G$-action in Lemma 4.6.
Proof.
Since the action $G$ on $\Sigma$ is the restriction of the action $G\curvearrowright\mathcal{G}$, for any $g\in G$ the left multiplication satisfies $L\circ\varphi_{g}=\varphi_{g}\circ L$. Thus, the $G$-equivariance of $\mathcal{J}_{0}$ follows from the $G$-equivariance of $\mathcal{J}$, that is $Ad_{g}^{*}\mathcal{J}_{0}=\mathcal{J}\circ\varphi_{g}\circ L=\mathcal{J}_{0}\varphi_{g}.$
Moreover, for $v\in\mathfrak{g}$, its fundamental vector field $v_{\mathcal{G}}|_{\Sigma}$ is actually a vector field on $\Sigma$. Hence, $L^{*}\imath_{v_{\mathcal{G}}}\omega=\imath_{v_{\mathcal{G}}}L^{*}\omega$. Since $\mathcal{J}$ is a cosymplectic moment map and $\xi(\langle J,v\rangle)=0$, we conclude
$d\langle\mathcal{J}_{0},v\rangle=L^{*}d\langle\mathcal{J},v\rangle=L^{*}(\imath_{v_{\mathcal{G}}}\omega+\xi\langle\mathcal{J},v\rangle\eta)=L^{*}\imath_{v_{\mathcal{G}}}\omega=\imath_{v_{\mathcal{G}}}L^{*}\omega.$
Under the above assumptions, the conditions of cosymplectic Hamiltonian action and symplectic Hamiltonian action are satisfied. Therefore, we can perform the structure reductions. These two reductions can be done simultaneously by observing that the diagram
[…]
commutes. The reduced symplectic groupoid $\mathcal{J}_{0}^{-1}(0)/G$ is constructed by considering the symplectic structure as in [23], which is compatible with the structural maps of the groupoid [17, Prop. 4.6]. As a direct consequence we get the next theorem.
Theorem 4.8.
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ a cosymplectic moment map satisfying the conditions of Theorem 3.4. Let $\Sigma$ the symplectic leaf containing $M$. Then
$L_{red}^{*}\omega_{red}=(L^{*}\omega)_{red},$
where $L:\Sigma\to\mathcal{G}$ is the inclusion and $L_{red}:\mathcal{J}_{0}^{-1}(0)/G\to\mathcal{J}^{-1}(0)/G$.
Proof.
Since $G$ acts freely and properly on $M$, the $G$-action on $\Sigma$ is also free and proper [17]. Thus, by a similar argument as in Remark 2.5, 0 is a regular value of $\mathcal{J}_{0}$. Moreover, $(\mathcal{J}_{0}^{-1}(0)/G,(L^{*}\omega)_{red})\rightrightarrows M/G$ is a symplectic groupoid [17]. The characterization of the reduced symplectic form on $\mathcal{J}_{0}^{-1}(0)/G$ is
$L_{0}^{*}(L^{*}\omega)=P_{0}^{*}(L^{*}\omega)_{red},$
where $L_{0}:\mathcal{J}_{0}^{-1}(0)\to\Sigma$ is the inclusion map and $P_{0}:\mathcal{J}_{0}^{-1}(0)\to\mathcal{J}_{0}^{-1}(0)/G$ is the quotient map. By the commutative in (11), we have
$L_{0}^{*}L^{*}\omega=L^{*}i^{*}\omega=L^{*}P^{*}\omega_{red}=P_{0}^{*}L_{red}^{*}\omega_{red}.$
Therefore, $P_{0}^{*}(L^{*}\omega)_{red}=P_{0}^{*}L_{red}^{*}\omega_{red}$. As usual, $P_{0}$ is a surjective submersion, so the desired result is obtained. ∎
4.3. Reduction of the multiplicative Chern class
In [15] the authors establish a relation between $S^{1}$-central extensions with vanishing multiplicative Chern class and cosymplectic groupoids. We use this identification and Theorem 3.4 to show that if the multiplicative Chern class vanishes, the same happens in the quotient.
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a Lie groupoid with a closed multiplicative 2-form, such that for any $x\in M$, $\text{rank}(\omega_{x})=2\dim(M)$. The pair $(\mathcal{G},\omega)$ is known as an oversymplectic groupoid [15, 7]. Furthermore, we suppose that $\mathcal{G}$ is orientable and proper. Assume that the foliation defined by $\ker\omega$ is simple, i.e., the leaf space $\mathcal{G}/\ker\omega$ is a smooth manifold and the map $\mathcal{G}\to\mathcal{G}/\ker\omega$ is a surjective submersion. Furthermore, assume that the corank of the foliation $\ker\omega$ is 1. Under these conditions, there exists a $S^{1}$-central extension
[…]
where $(\Sigma_{1},\omega_{1})$ is a symplectic groupoid and $q^{*}\omega_{1}=\omega$.
Moreover, $\mathcal{G}$ becomes a $S^{1}$-principal bundle. Let $\partial_{\theta}$ be the infinitesimal generator of the $S^{1}$- action on $\mathcal{G}$. Because $\mathcal{G}$ is a proper groupoid there exists a multiplicative form $\eta\in\Omega^{1}(\mathcal{G})$ such that $\imath_{\partial_{\theta}}\eta=1$. The form $\eta$ is related to the multiplicative Ehresmann connection, and the curvature of this connection is $\Omega=d\eta$. For more details on this construction the reader can consult [15, 16].
The 2-form $\Omega$ is multiplicative and $q$-basic i.e., there exists a form $\Omega_{1}\in\Omega_{cl}(\Sigma_{1})$ such that $q^{*}\Omega_{1}=\Omega$. The multiplicative Chern class is defined as the cohomology class associated with $\Omega_{1}$ (in the multiplicative de Rham cohomology). One of the main results in [15, Thm. 4.12] states that there exists a multiplicative 1-form $\eta$ such that $(\mathcal{G},\omega,\eta)$ is a cosymplectic groupoid if and only if the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
For a cosymplectic groupoid the foliation defined by the kernel of the 2-form is simple. On the other hand, if $\mathcal{G}$ is a proper groupoid, and there is an $G$-action on $\mathcal{G}$ and a cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ as in Theorem 3.4, the groupoid $\mathcal{G}_{red}\rightrightarrows M/G$ may be a non-proper groupoid. However, with additional conditions the reduced groupoid is proper.
Lemma 4.9.
[…]
Proof.
Since the action $G$ on $\mathcal{H}$ is free and proper [17], the quotients $\mathcal{H}/G$ and $M/G$ are smooth manifolds. Provided that the structural maps of $\mathcal{H}\rightrightarrows M$ are $G$-equivariant, then they descend to $\mathcal{H}/G$. Now, we show that under these hypotheses $\mathcal{H}$ is a proper groupoid. To show that $(s,t):\mathcal{H}\to M\times M$ is a proper map, we consider a sequence of arrows $y_{n}\xleftarrow{g_{n}}x_{n}\in\mathcal{H}$ such that $(x_{n},y_{n})\to(x,y)\in M\times M$, and we show that $g_{n}$ admits a convergent subsequence in $\mathcal{H}$ (see [12, 4]). Since $\mathcal{G}$ is a proper groupoid, there exists a subsequence $g_{n_{k}}\to g\in\mathcal{G}$. Because $\mathcal{H}\subset\mathcal{G}$ is closed, then $g\in\mathcal{H}$.
Let $[y_{n}]\xleftarrow{[g_{n}]}[x_{n}]\in\mathcal{H}/G$ such that $([x_{n}],[y_{n}])\to([x],[y])\in M/G\times M/G$. Thus, it is enough to show that $[g_{n}]$ admits a convergent subsequence in $\mathcal{H}/G$. Since $(p,p):M\times M\to M/G\times M/G$ and $(s,t):\mathcal{H}\to M\times M$ are proper maps, there exists a subsequence $y_{n}\xleftarrow{g_{n_{k}}}x_{n}$ converging to $g\in\mathcal{H}$. By the continuity of the map $\mathcal{H}\to\mathcal{H}/G$, we conclude that $[g_{n_{k}}]\to[g]\in\mathcal{H}/G$. ∎
In Lemma 4.9, considering the group $G$ to be compact, the condition that the map $M\to M/G$ be proper is changed to be a closed map. Next, we state a result which follows directly from Theorem 3.4 and the characterization between cosymplectic groupoids and $S^{1}$-central extension with vanishing multiplicative Chern class presented in [15, Thm. 4.12].
Proposition 4.10.
Let $G$ be a Lie group acting on a cosymplectic groupoid $(\mathcal{G},\omega,\eta)\rightrightarrows M$, and $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ a cosymplectic moment map satisfying the conditions of Theorem 3.4. If $p:M\to M/G$ is a proper map, then for $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
Proof.
By using Theorem 3.4 we obtain a cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$. Since $M\to M/G$ was assumed to be a proper map, by Lemma 4.9, $\mathcal{G}_{red}$ is proper. Consequently, by using the result in [15, Thm. 4.12], we conclude that the corresponding $S^{1}$-central extension of $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ has vanishing multiplicative Chern class. ∎
If we start with a corank 1, orientable, proper oversymplectic groupoid, with simple foliation $\ker\omega$, such that the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class, the following theorem gives a way to produce new groupoids where the multiplicative Chern class vanishes.
Theorem 4.11.
Let $(\mathcal{G},\omega)\rightrightarrows M$ be a corank 1, orientable, proper oversymplectic groupoid, with simple foliation $\ker\omega$, such that $H^{1}(\mathcal{G})=0$. Let $G$ be a compact Lie group acting on $\mathcal{G}$ by Lie groupoid automorphisms preserving $\omega$. The following is true:
  1. (1)
    If the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class, then there exists a cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ for a cosymplectic groupoid structure on $\mathcal{G}$, such that $\xi(\langle\mathcal{J},v\rangle)=0$ for all $v\in\mathfrak{g}$.
  2. (2)
    If, in addition, $\mathcal{J}$ is a groupoid morphism, $G$ acts freely on $M$, and $M\to M/G$ is a closed map, then for $(\mathcal{G}_{red},\omega_{red},\eta_{red})\rightrightarrows M/G$ the corresponding $S^{1}$-central extension has vanishing multiplicative Chern class.
Proof.
To prove item (1), the condition on the multiplicative Chern class defines a cosymplectic structure $(\mathcal{G},\omega,\eta)$ [15, Thm. 4.12]. Because $G$ is compact, we consider the average
[…]
where $d\Lambda$ is the Haar measure of the compact group $G$. From Haar measure invariance it follows that $\bar{\eta}$ is a $G$-invariant, closed, multiplicative 1-form in $\mathcal{G}$.
We claim that $(\mathcal{G},\omega,\bar{\eta})$ is a cosymplectic manifold. In fact, note that the Reeb vector field $\xi$ of $(\mathcal{G},\omega,\eta)$ satisfies $\imath_{\xi}\omega=0$ and $\imath_{\xi}\bar{\eta}\neq 0$. Thus, if the map $\flat:\mathfrak{X}(\mathcal{G})\to\Omega^{1}(\mathcal{G})$ defined in (1) is not injective, we can consider $X\neq 0$ in the kernel of $\flat$. Note that $0=\imath_{\xi}\flat(X)=(\imath_{X}\bar{\eta})(\imath_{\xi}\bar{\eta})$. Hence, $\bar{\eta}$ vanishes on $X$, and $\imath_{X}\omega=\flat(X)=0$. Thus, $X=f\xi$ for some non-vanishing smooth function $f$. However, $\imath_{X}\bar{\eta}=f\imath_{\xi}\bar{\eta}$, which is a contradiction. Consequently, $\flat$ is an isomorphism. We conclude this claim, by considering the normalization $\bar{\xi}$ of $\xi$, such that $\imath_{\bar{\xi}}\bar{\eta}=1$, and Remark 2.2.
Since $\bar{\eta}$ is invariant by the action, we get that $d(\imath_{v_{\mathcal{G}}}\bar{\eta})=\mathcal{L}_{v_{\mathcal{G}}}\bar{\eta}=0$. Because $\epsilon^{*}\bar{\eta}=0$, then $0=\imath_{v_{M}}\epsilon^{*}\bar{\eta}|_{x}=\imath_{v_{\mathcal{G}}}\bar{\eta}|_{\epsilon(x)}$ for $x\in M$. Hence, $\imath_{v_{\mathcal{G}}}\bar{\eta}=0$ for all $v\in\mathfrak{g}$. The existence of the cosymplectic moment map $\mathcal{J}:\mathcal{G}\to\mathfrak{g}^{*}$ follows form Proposition 2.6.
To prove item (2), we consider the reduced cosymplectic groupoid $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ obtained from Theorem 3.4. By using Lemma 4.9, $\mathcal{G}_{red}$ is proper. From [15, Thm. 4.12], we conclude that the corresponding $S^{1}$-central extension of $(\mathcal{G}_{red},\omega_{red},\bar{\eta}_{red})\rightrightarrows M/G$ has vanishing multiplicative Chern class. ∎
5. Further direction of study
  • Recall that in the cosymplectic reduction Theorem 2.4 it is assumed that $v_{Q}$ is the Hamiltonian vector field of $\langle J,v\rangle$ and $\xi\langle J,v\rangle=0$, this in particular implies that $\mathcal{L}_{v_{Q}}\omega=0$ and $\mathcal{L}_{v_{Q}}\eta=0$. However, we can suppose conformal symmetries, which leads us to consider the conditions
    $\mathcal{L}_{v_{Q}}\omega=-d(\xi\langle J,u\rangle)\wedge\eta\text{\ and\ }\mathcal{L}_{v_{Q}}\eta=f_{v}\eta$
    for some non-vanishing function $f_{v}$ for each $v\in\mathfrak{g}$. In this context, the reduction theorems are different, and would state some other relation with respect to the cosymplectic groupoids reduction.
  • In [15, Prop. 3.3] the authors state an infinitesimal identification with a Lie bialgebroid of the form $(T^{*}M\oplus\mathbb{R},TM\oplus\mathbb{R})$. However it is also known the existence of generalized Lie bialgebroids [21] and triangular generalized Lie bialgebroids as in [20]. The reduction results of this paper can be used to reduce such bialgebroids [21]. They can also be used to compare other geometric terms, such as the modular class in the reduction presented in [20].
  • Integration of Poisson structures can also be understand via Poisson sigma models [10]. As we mention previously, if the groupoid $\mathcal{G}$ is cosymplectic, it has a Poisson bracket. Furthermore, the units $M$ also has a Poisson structure and a Lie algebroid structure. Results in [15] and in these notes can be extended to the integration theory given in [10] via sigma models.
  • The construction of the moment map in Section 4.1 comes from the ideas in [6], where there is an action of a Lie group on an algebroid which lifts an action on the basis. By the Lie’s second theorem, this action is integrated to an action of a Lie group on a Lie grouproid. It would be valuable to see if the ideas worked out here can be generalized by considering the action of a 2-group on a cosymplectic groupoid and its infinitesimal counterpart.
Statements and Declarations: The authors have no conflict of interest to declare that are relevant to this article.
Data Availability Statements: Data sharing not applicable to this article as no datasets were generated or analysed during the current study.